Order posts by limited to posts

12 Mar 12:01:00
12 Mar 09:50:23
TalkTalk have a a fibre outage affecting EoFTTC lines in the East Midlands and East Anglia area. Lines will be in sync, but will not have any further connectivity. TalkTalk are aware and have engineers investigating. They have their own status post here: http://solutions.opal.co.uk/network-status-report.php?reportid=6291
12 Mar 10:18:51
Not further updates as yet.
12 Mar 10:20:37
Update from TT: Our NOC and Network Support Engineers have confirmed that corrective work to restore services to all impacted exchanges remains on-going.
12 Mar 11:44:30
Update from TalkTalk: "Our NOC Engineers have confirmed that all network alarms have cleared and normal working service is now fully restored to all impacted exchanges and customers. There are continued reports of some B2B circuits continuing to experience issues and our IP Engineers are currently investigating these reports. Incident Management will continue to liaise with our NOC Engineers and further updates will be provided upon receipt."
12 Mar 14:20:48
Service has been restored.
Started 12 Mar 00:10:00
Closed 12 Mar 12:01:00

6 Mar 18:14:21
6 Mar 16:05:00
We have identified that Ethernet customers are not seeing archived graphs for last few days and the cause is part of the recent router upgrade code. We plan to carry out a further upgrade this evening which should have little or no disruption, and should rectify this issue.
6 Mar 16:53:29
There may be a blip on some TalkTalk lines during this upgrade.
Resolution Work completed for now - seems to have gone as planned. There was a small TalkTalk line blip as expected.
Started 6 Mar 17:00:00
Closed 6 Mar 18:14:21
Previously expected 6 Mar 19:00:00

24 Feb 2014 16:00:00
24 Feb 2014 16:01:15
Some work was done today (following issues at the weekend). This work is on part of a redundant pair of routers, and so should have no impact. The normal redundancy aspects have worked, with changes to routing and VRRP, but for some unknown reason there are issues with BGP announcements internally which are causing blips on external connectivity to some Ethernet customers. The work has been completed but we are now investigating how this has had an impact on services so they can be avoided in future.
Started 24 Feb 2014 15:45:00
Closed 24 Feb 2014 16:00:00
Previously expected 24 Feb 2014 16:00:00

02 Sep 2011 09:41:55
02 Sep 2011 09:43:14

The usage graphs were disrupted yesterday due to various work on core routers, and will be reset today starting shortly. This is to fix a problem where the ping monitoring was not working correctly. Service should otherwise be unaffected by this change.

Started 02 Sep 2011
Closed 02 Sep 2011 09:41:55

30 Aug 2011 22:30:00
30 Aug 2011 13:34:53

There appears to be a major issue with fibre links in to our node in Maidenhead. This is also affecting out office which is why there was a delay posting this.

Staff are on the way to site now to investigate.

30 Aug 2011 13:43:01

Staff are on the way. As ever, backup equipment was not set up right in our offices which had led to a delay updating customers on this. Incoming calls are being handle by very few staff at present which is also causing delays taking calls. We expect staff to be back on it in a few minutes.

30 Aug 2011 14:02:08

From what we can see most things in the data centre are working, e.g phones. But the fact we are unable to call the data centre suggests they have something major going on. More to follow. Staff still not managed to get on site yet.

30 Aug 2011 14:10:31

Ok still no proper news, but the data centre is apparently swamped with openreach, suggesting something major, but also suggesting that thy are on the case.

30 Aug 2011 14:16:50

We can confirm that all the fibre links to our rack in Maidenhead are showing alarms at present, but that power and everything else is fine. BT vans are on site and they have a big hole in the ground, but the BT engineers claim nothing wrong and they did not touch anything.

We are reporting multiple fibre failures to BT now.

30 Aug 2011 14:18:05

Oh, and no clue why data cent were not answering the phone. We'll worry about that later.

30 Aug 2011 14:57:13

We have confirmation that other customers in the data centre have also lost fibre links. BT are on the case but no ETA yet.

30 Aug 2011 15:40:21

Ok. Clarifying rumours here. BT vans on site, apparently on unrelated things. Hole in ground is near site and some sort of construction work. May be unrelated. Maybe not. We are waiting up updates from BT now.

30 Aug 2011 16:25:02

It appears that BT had internal system problems meaning they have not actually logged the fault internally until just before 4pm!!

30 Aug 2011 16:30:10

Whilst delays in BT are frustrating, the fact this appears to be a fibre break with other people affected means they should already be working on it. We hope. We are still awaiting an update.

30 Aug 2011 16:59:50

Further checks confirm that our link from our office to Maidenhead is only half broken, which is surprising. I.e. one of the two fibres is apparently working. This suggests a damaged rather than cut fibre.

30 Aug 2011 17:53:00

So far, all we have, is that BT are running diagnostics. No more information yet.

30 Aug 2011 18:05:53

Now waiting for a call from the Openreach engineer who is running diagnostics...

30 Aug 2011 18:37:00

There are two Openreach engineers on site investigating. They have said that they're investigating someone elses's fault on site so sounds like they might have found a common problem.

30 Aug 2011 18:53:47

BT engineer turned up on site to find two other engineers already working on it. So it is in hand. No ETA.

30 Aug 2011 19:55:10

BT found what looks to be a physical fibre fault (so a break somewhere - notes added at 19:24) and confirmed at 19:37 that they are still working on it.

30 Aug 2011 21:03:42

BT are still working on the break. No ETA yet.

30 Aug 2011 21:51:05

Still no news. BT should call back, hopefully with more information, soon.

30 Aug 2011 21:52:25

Whilst we expect BT to sort this over night, we may not have status updates until early morning. All systems are set up to come back on line as soon as BT fix the fibre.

30 Aug 2011 22:18:03

BT are trying to put in a temporary fix to the problem. No details as to the root cause, and no time scale for a fix yet, but they've said that if the temporary fix fails they'll have to replace the broken fibre. If that happens it is unlikely to be fixed tonight.

30 Aug 2011 22:42:26

Most affected customers seem to have come back online in the last few minutes.

30 Aug 2011 22:54:46

BT have confirmed that this is a "permanent fix", by which they mean that any further disruption should be a part of planned engineering works.

31 Aug 2011 08:08:58

Our links for Ethernet services have been restored and our office is back on line. We're waiting for a full explanation from BT.

31 Aug 2011 13:25:09

More details: http://aa.net.uk/news-2011-fibre-break.html


BT fixed broken fibre.

Started 30 Aug 2011 13:14:00
Closed 30 Aug 2011 22:30:00

27 Aug 2010 17:44:00
27 Aug 2010 17:54:41

At just before 17:00 today a minor routing announcement change was made. This was a very minor change and was not run as a proper planned work because it was so minor, or so we thought. Even so the change was not done in the middle of the day.

At 17:36 we were advised by MSO text that customers on Ethernet links were seeing problems. By 17:44 we had identified the cause of the problem and resolved it.

Technical details: The problem is in fact related to exhaustion of IPv4 space. We are involved in arrangements to swap IP blocks around between ISPs to free up some additional space, and this meant fragmenting one of our larger IP blocks. In principle this is simple, we announce the smaller blocks and drop the larger block announcement later. Unfortunately one of the smaller blocks is routed to Ethernet connected customers and was included in the list annonced at Telehouse in London.

The impact was that some of the Internet was not accessable to Ethernet customers, depending on routing. This included all of our ADSL lines and many services peering on LINX in London which could not see Ethernet lines. Unfortunately, as a partial failure it was not picked up by any of our automated monitoring which could still see Ethernet customers.

This was human error, and resolved within 8 minutes of being reported on a bank holiday weekend out of hours.

We will consider ways to manage this better - it is a trade off as we could have put more in automated control which would have stopped this specific error but could lead to more radical errors being possible with a lot less effort. We're reviewing procedures for checking changes like this before hand in future.

Sorry for the inconvenience.

Started 27 Aug 2010 17:00:00
Closed 27 Aug 2010 17:44:00